Cost Behavior, Resource Usage, Excess Capacity

Rolertyme Company manufactures roller skates. With the exception of the rollers, all parts of the skates are produced internally. Neeta Booth, president of Rolertyme, has decided to make the rollers instead of buying them from external suppliers. The company needs 100,000 sets per year (currently it pays $1.90 per set of rollers).

The rollers can be produced using an available area within the plant. However, equipment for production of the rollers would need to be leased ($30,000 per year lease payment). Additionally, it would cost $0.50 per machine hour for power, oil, and other operating expenses. The equipment will provide 60,000 machine hours per year. Direct material costs will average $0.75 per set, and direct labor will average $0.25 per set. Since only one type of roller would be produced, no additional demands would be made on the setup activity. Other overhead activities (besides machining and setups), however, would be affected. The company’s cost management system provides the following information about the current status of the overhead activities that would be affected. (The supply and demand figures do not include the effect of roller production on these activities.) The lumpy quantity indicates how much capacity must be purchased should any expansion of activity supply be needed. The purchase price is the cost of acquiring the capacity represented by the lumpy quantity. This price also represents the cost of current spending on existing activity supply (for each block of activity).

Activity Price Cost Driver Supply Usage Lumpy
Quantity
Purchase
Purchasing Orders 25,000 23,000 5,000 $25,000
Inspection Hours 10,000 9,000 2,000 30,000
Materials handling Moves 4,500 4,300 500 15,000

Production of rollers would place the following demands on the overhead activities:

Activity Resource Demands
Machining 50,000 machine hours
Purchasing 2,000 purchase orders (associated with raw materials used to make the rollers)
Inspection 750 inspection hours
Materials handling 500 moves

Producing the rollers also means that the purchase of outside rollers will cease. Thus, purchase orders associated with the outside acquisition of rollers will drop by 5,000. Similarly, the moves for the handling of incoming orders will decrease by 200. The company has not inspected the rollers purchased from outside suppliers.

Required:

1. Using the questions below, classify all resources associated with the production of rollers as flexible resources and committed resources and as a short- or long-term commitment.

a. Direct materials, direct labor and machine operating costs would be classified as:

flexible resources

b. Machining would be classified as:

long-term committed resources

c. Purchasing, inspection and materials handling would be classified as:

short-term committed resources

2. Calculate the total annual resource spending (for all activities except for setups) that the company will incur after production of the rollers begins. Break this cost into fixed and variable activity costs. In calculating these figures, assume that the company will spend no more than necessary.

Fixed Cost $
Variable Cost $
Total Cost $

What is the effect on resource spending caused by production of the rollers?

Decrease   = $

3. Refer to Requirement 2. For each activity, break down the cost of activity supplied into the cost of activity output and the cost of unused activity.

Activity Cost of Activity Supplied Cost of Activity Used Cost of Unused Activity
Machining $ $ $
Purchasing
Inspection
Materials handling
Answer #1
  1. Flexible resources:

Direct materials, Direct Labour and Machine operating costs

Committed resources:

                       Short term: Purchasing, inspection and materials handling

                       Long term: Machining

                Both long term and short term are considered as fixed activity costs.

  1. Total Annual Resource Spending:

Activity

Fixed costs

Variable costs

Total costs

Material usage

   -

75000

75000

Labour

-

25000

25000

Machining

30000

25000

55000

Purchasing

100000

-

100000

Inspection

150000

-

150000

Material Handling

150000

150000

Total

430000

125000

555000

Explanation:

Fixed costs:

Machining: Lease payment of $30,000 is taken as machining fixed cost.

Variable costs:

Material usage = $0.75 * 100000 = $75,000

Labour = $0.25 * 100000 = $25,000

Machining = $0.5 * 50000 = $25,000

Purchasing for required supply = 23000 – 5000 + 2000 = 20000

The resources are going to be purchased in the units of 5000, the supply can be reduced from 25000 to 20000.

Inspection: This resource is purchased in blocks of 2000. Demand = 9000 + 750 = 9750.

Therefore, supply must be equal to 10000. Spending = 10000/2000 * 30000 = $150000

Material handling: This resource is purchased in blocks of 500.

Demand = 4300 + 500 – 200 = 4600.

Supply must be 5000.

Resource spending = 5000/500 * 15000 = $150000

Effect on resource spending of decision to produce rollers:

Material

(75000)

Labour

(25000)

Machining

(55000)

Purchasing

25000

Inspection

0

Material handling

(15000)

Outside Purchase

190000

Total decrease

$45,000

Supply drops from 25000 to 20000 orders. Hence the saving is $25000

Activity Inspection is 10000 hrs, so no change in spending is required.

Activity material handling increases by 500 moves, spending increases by $15000

No need to do outside purchases therefore, savings = 100000 * $1.90 = $190,000

Activity

Cost of Activity Supplied ($)

Cost of activity used ($)

Cost of unused activity ($)

Machining

30000

25000

5000

Purchasing

100000

100000

0

Inspection

150000

146250

3750

Material handling

150000

138000

12000

Cost of unused activity = Fixed activity rate * unused capacity

Machining = ($30000 / 60000) * 10000 = $5000

Purchasing = ($100000 / 20000) * 0 = $0

Inspection = ($150000 / 10000) * 250 = $3750

Material handling = ($150000 / 5000) * 400 = $12,000

Cost of activity used = Cost of activity supplied – Cost of unused activity